1. You have in your possession an acid made up of 68.1% oxygen, 29.8% nitrogen, and the rest hydrogen. Assuming that the molecular formula is the same as the empirical formula, determine the formula of your acid.

2. Name the acid
3. Draw the lewis structure
4. Give the oxidation numbers of each element in acid.
5. Does your acid contribute to acid rain? If so explain how, using balanced chemical reactions and listening sources of your pollutant.
1
Expert's answer
2012-10-19T09:00:48-0400
1 )68.1/16 : 29.8/14 : 2.1/1 = 4.25 :2.13 : 2.1=2:1:1 O2NH

2) it is HNO2
Nitrous acid

3) H-O-N=O H:O:N::O

4) H -- 1 , O -- 2 , N -- 3


5) Yes N2 + O2 = 2NO in the rain during lightning

2NO + O2=2NO2
2NO2 + H2O = HNO2 + HNO3